line PQ parallel to line AB.Slope of line AB is 1/2 . Y-intercept of line PQ is 3 find X- intercept.
Answer by Boreal(15235)

PQ has a slope of 1/2, since parallel lines have the same slope
The line has equation of y=(1/2)x+3
the x-intercept has y=0, so 0=(1/2)x+3
(1/2)x=-3
x=-6
(-6, 0)